mathslover (mathslover):

similar one but confused: \[\large{\frac{a}{b}=\frac{b}{c}=\frac{c}{d}}\] prove that: \(\large{\frac{a}{d}=}\) \(\large{\sqrt{\frac{a^5+b^2c^2+a^3c^2}{b^4c+d^4+b^2cd^2}}}\)

mathslover (mathslover):

\[\large{\sqrt{\frac{a^5+b^2c^2+a^3c^2}{b^4c+d^4+b^2cd^2}}=\frac{a}{d}}\]

mathslover (mathslover):

tried to put b = \(\large{\sqrt{ac}}\) and c=\(\large{\sqrt{ad}}\)

mathslover (mathslover):

sorry : \[\large{c=\sqrt{bd}}\]

mathslover (mathslover):

@asnaseer ...

OpenStudy (anonymous):

me confused too...

mathslover (mathslover):

yes

OpenStudy (anonymous):

maybe there is a easier way but u can factor out \(a^2\) in num and \(d^2\) in denum Under the radical .... then firstly replace \(d^2\)'s with \(b^2c^2/a^2\) and then replace all \(b^2\)'s with \(ac\)

OpenStudy (anonymous):

wait...there is an easier way...

OpenStudy (anonymous):